OpenStudy (anonymous):

The diagonals of parallelogram ABCD intersect at point E. If DE=2X+2,BE=3X-8 ,CE=4y , and AC=32, solve for x. A. x = 2 B. x = 10 C. x = 15 D. x = 22

OpenStudy (anonymous):

Solve for x & y.

OpenStudy (anonymous):

the diagonals bisect eachother. 2x+2 = 3x -8 and 4y = 1/2 (32)

OpenStudy (anonymous):

with geometry a lot of it is just memorizing those kind of relationships to help you solve using algebra

OpenStudy (anonymous):

4/ by 32= 1/8 ?

OpenStudy (anonymous):

thats .5

OpenStudy (anonymous):

y=8 and x=?

OpenStudy (anonymous):

4y = 16 ....

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yep! with the other one you need to get the x's on the same side, and solve

OpenStudy (anonymous):

6=-2 ?

OpenStudy (anonymous):

x=2 ?

OpenStudy (anonymous):



OpenStudy (anonymous):

2x+3x=2-8

OpenStudy (anonymous):

start by getting variables on same side. 2x + 2 - 2x = 3x - 8 - 2x 2 = x -8

OpenStudy (anonymous):

x+2=-1x-8

OpenStudy (anonymous):

close, but you have to change the sign in order to get it across the equation

OpenStudy (anonymous):

???

OpenStudy (anonymous):

how

OpenStudy (anonymous):

do you see how i got 2 = x - 8?

OpenStudy (anonymous):

once you subtract 2x from both sides, it's gone on the left, and 3x-2x = x on the right

OpenStudy (anonymous):

so subtract 2 from 3x-2=x ??

OpenStudy (anonymous):

3x=-2x

OpenStudy (anonymous):

Does x=2 ?

OpenStudy (anonymous):

sorry, the 3x-2x part is the right side ONLY. on the left side we did 2x + 2 -2x = 2

OpenStudy (anonymous):

x - 8 = 2

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Yep! when doing the alebra, write it out line by line and you'll see what i mean. It's hard to explain on here

OpenStudy (anonymous):

haha do you know how to use the draw button? x) it makes things simpler.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

lol i was looking for it but dont have my stylus anyway, no prob